OpenStudy (anonymous):

The function R(x)=-0.001x@Sup{2} + 3x describes the money that the local sandwich store takes in each week from the sale of x roast beef sandwiches. Use this revenue function and the cost function (.45x + 535) to write the store's weekly profit function, P. (Hint: The profit function is the difference between revenue and cost functions.)

OpenStudy (helder_edwin):

what is "@Sup"

OpenStudy (anonymous):

"squared" in superscript. It didn't translate over well. In other words, it could be read -0.001x"squared"

OpenStudy (helder_edwin):

so it is \[ \large R(x)=-0.001x^2+3x \] and \[ \large C(x)=0.45x+535 \]

OpenStudy (helder_edwin):

\[ \large P(x)=R(x)-C(x) \] do u know calculus?

OpenStudy (anonymous):

Unfortunately no. This is for algebra studies.

OpenStudy (helder_edwin):

ok. then u must know how to complete the square

OpenStudy (anonymous):

i think all you need to do now is to just simplify the expression R(x) - C(x)

OpenStudy (anonymous):

The answer I came up with seemed like exactly that - that it needed to be simplified, but I don't see how to simplify it past this: f = 2.55 + -535x-1 + -0.001x

OpenStudy (anonymous):

the question just asks for the profit function...

OpenStudy (anonymous):

Right. I should have made it equal to p(x)

OpenStudy (helder_edwin):

yes. u r right @dpaInc . i assumed it was asking for the maximum profit.

OpenStudy (anonymous):

yeah... i had to read the question several times myself... i misread the problem too...:)

OpenStudy (helder_edwin):

ok @bgarrett \[ \large P(x)=R(x)-C(x)=-0.001x^2+3x-(0.45x+535) \]

OpenStudy (helder_edwin):

just simplify like terms as dpaInc said

OpenStudy (anonymous):

Things guys. That works. I plugged it into the online exam. I appreciate it!

OpenStudy (helder_edwin):

u r w

OpenStudy (anonymous):

:) Thanks. Let me add a few typos to the already confusing world of mathematics.